OpenStudy (anonymous):

Please Help v.v 9p + 8 = 10p + 7 A) Open; there is a variable. B) True; the expressions are the same for all values. C) False; the expressions are never the same.

OpenStudy (anonymous):

I got B but not sure if I am right..

OpenStudy (anonymous):

Oh. :/

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

It's A and if I am wrong I am sorry.

OpenStudy (anonymous):

Okay ._.

OpenStudy (anonymous):

Have you completed this test? for 9th grade :I

OpenStudy (anonymous):

I am in 11th haven' done this stuff in a while.

OpenStudy (anonymous):

Haven't*

OpenStudy (anonymous):

Wow v.v

OpenStudy (anonymous):

Do you know whose in the 9th grade that can help me more because I have more where that came from v.v

OpenStudy (anonymous):

Did I get it right?

OpenStudy (anonymous):

I don't know yet still got questions left over v.v

OpenStudy (anonymous):

Ah

OpenStudy (anonymous):

Shoot

OpenStudy (anonymous):

._.

OpenStudy (anonymous):

As in give me a few I will see what I come up with

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Which Ordered pair is a solution of equation y = x - 4 A) 2,6 B) 6,2 C) -2,6 D-6,2

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

You sure?

OpenStudy (anonymous):

im backwards its B

OpenStudy (anonymous):

noits A

OpenStudy (anonymous):

Because 6-4=2

OpenStudy (anonymous):

Oh :o

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Which Ordered pair is a solution equation y = 3x? A) -2,-9 B) -8,-18 C) -8, -3 D) -10,-30

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Which ordered pair is a solution of the equation y = -7 + 2 A) 1,2 B) 8,-54 C)5,-35 D)1,7

OpenStudy (anonymous):

D

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

o.o Okay.

OpenStudy (anonymous):

yeah D

OpenStudy (anonymous):

I'm going to take a snap shot of questions :/

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

http://gyazo.com/30d86ef6b3d9569454e71fb9967c349d

OpenStudy (anonymous):

Cant view gyayzo

OpenStudy (anonymous):

Gyazo*

OpenStudy (anonymous):

Really ._.

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

I can view lightshot

OpenStudy (anonymous):

Hmm. .

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

http://prntscr.com/4nq7md